1-10) Equation:
1) a (amplitude) v) period ii) b (frequency) iii) c (phase shift) iv) d (vertical shift)
Answer #1

Equation a sin(4x) +2 (i) amplitude = 2 81 (1) frequency = 21 Tay so w= 27 = 7 (iii) phase shiftzo (iv) vertical shift=2 (V)